OpenStudy (anonymous):

A plane is missing and is presumed to have equal probabilityof going down in any of three regions. If a plane is actually downin region i, let 1-αi denote the probability thatthe plane will be found upon search of the ith region, i=1,2,3.What is the conditional probability that the plane is in a) region 1, given that the search of region 1 wasunsuccessful b) region 2, given that the search of region 1 wasunsuccessful c) region 3, given that the search of region 2 wasunsuccessful

OpenStudy (anonymous):

oh i see it says the probability is \(1-\alpha_i\) is that right?

OpenStudy (anonymous):

Yes that is right

OpenStudy (anonymous):

I know I have to use the Bayes Theorem, but I'm not understanding it at all when i try to apply it.

OpenStudy (anonymous):

If I can get through A I think I'll understand the rest

OpenStudy (anonymous):

ok now i have to think

OpenStudy (anonymous):

dont think you need baye's for the first one if we put A is the event the plane went down in region 1, and B is the event the search was unsuccessful, then you are looking for \[P(A|B)=\frac{P(A\cap B)}{P(B)}\] the denominator is \(\alpha_1\) and the numerator is \(\frac{1}{3}\times (\alpha_1)\)

OpenStudy (anonymous):

maybe i am messing this up, let me think for another moment

OpenStudy (anonymous):

except isn't \[(1-\alpha _{i})\] the probability of finding the plane GIVEN that it went down in the region? (P(B|A))?

OpenStudy (anonymous):

give me a minute, if you have it

OpenStudy (anonymous):

yeah to your answer, but the given part is "the search is unsuccessful"

OpenStudy (anonymous):

the language is confusing, and also the fact that the condition is not "the search is unsuccessful' but "the search in region 1 is unsuccessful'

OpenStudy (anonymous):

I think the problem is that I don't know what P(B) is, which is where we'd apply Bayes Theorem/ Law of Total Probability.

OpenStudy (anonymous):

ok i think i got it at least for the first one

OpenStudy (anonymous):

lets put A as the even the plane went down in region 1, B as the event the search is unsuccessful in region 1

OpenStudy (anonymous):

then \[P(A|B)\] is what you are looking for

OpenStudy (anonymous):

bayes gives \[P(A|B)=\frac{P(B|A)P(A)}{P(B|A)P(A)+P(B|A^c)P(A^c)}\]

OpenStudy (anonymous):

if the plane is down in region 1, then the probability it is found is \(1-\alpha_1\) so the probability it is not found is \(\alpha_1\) i.e. \[P(B|A)=\alpha_1\]

OpenStudy (anonymous):

and of course \(P(A)=\frac{1}{3}\)

OpenStudy (anonymous):

\[P(B|A^c)\] is the probability that the search is region 1 is unsuccessful given that the plane is not in region 1, and that is 1

OpenStudy (anonymous):

and \(P(A^c)=\frac{2}{3}\)

OpenStudy (anonymous):

okay that all makes sense so far

OpenStudy (anonymous):

we put it together and get \[\frac{\alpha_1\times \frac{1}{3}}{\alpha_1\times \frac{1}{3}+\frac{2}{3}}\] multiply top and bottom by 3 and get \[\frac{\alpha_1}{\alpha_1+2}\]

OpenStudy (anonymous):

i would not bet $20 that this is right, but i would bet $12

OpenStudy (anonymous):

Well this makes sense with my (very limited) understanding of the material

OpenStudy (anonymous):

what about part B then?

OpenStudy (anonymous):

ok i think this one is a bit harder, but doable

OpenStudy (anonymous):

my professor did say something about it being a bit of a trick question

OpenStudy (anonymous):

this time lets put A and the plane is in region 2, and B as the search in region 1 is unsuccessful

OpenStudy (anonymous):

we want \[P(A|B)\] and again it is \[\frac{P(B|A)P(A)}{P(B|A)P(A)+P(B|A^c)P(A^c)}\]

OpenStudy (anonymous):

\[P(B|A)\] is the probability that the search in region 1 is unsuccessful given that the plane is in region 2, and that is 1

OpenStudy (anonymous):

the part that is not obvious is \(P(B|A^c)\) and for that we need two pieces. this is the probability that the search in region 1 is unsuccessful given that the plane is not in region 2

OpenStudy (anonymous):

for our denominator what we really need is the probability of B, i.e. we want the probability that the search in region 1 is unsuccessful so we need to consider all three possibilities the plane is in region 1, the probability it is unsuccessful is \(\alpha_1\) the plane is in region 2, the probability is 1 the plane is in region 3, the probability is also 1 multiplying that by the probability that the plane is in each region give \[P(B)=\frac{1}{3}\alpha_1+\frac{1}{3}+\frac{1}{3}\]

OpenStudy (anonymous):

you will notice that this is exactly the same denominator we had before, i.e. we already computed \(P(B)\) so this was kind of a waste of time

OpenStudy (anonymous):

so for number 2, the answer is \[\frac{\frac{1}{3}}{\frac{1}{3}\alpha_1+\frac{2}{3}}\] \[=\frac{1}{\alpha_a+2}\]

OpenStudy (anonymous):

why do we need P(B)?

OpenStudy (anonymous):

because bayes formula is really nothing more than a method of computing \[P(A|B)=\frac{P(A\cap B)}{P(B)}\]

OpenStudy (anonymous):

the point is that sometimes you do not know \(P(B)\) directly so you have to find it somehow

OpenStudy (anonymous):

what we did in problem 1 was find the same thing, but we didn't know \(P(B)\) directly , not did we know \(P(A\cap B)\) but it is always the case that \(P(A\cap B)=P(B|A)P(A)\)

OpenStudy (anonymous):

in other words bayes formula is just a trick of algebra, but it allows you do change the conditioning from A given B to B given A if you know the second, then you can compute the first

OpenStudy (anonymous):

by the way, now that we know if B is the probability that the search in region 1 is unsuccessful, then it is pretty clear that the probability that the search in region 2 is unsuccessful is \[\frac{1}{3}\alpha_2+\frac{2}{3}\]

OpenStudy (anonymous):

btw we should have computed these first, it would have made life easier let me know if i have completely lost you

OpenStudy (anonymous):

let me think for a moment

OpenStudy (anonymous):

ok we could have even used a baby tree diagram to do this

OpenStudy (anonymous):

okay I think I got it. So this time instead of changing the bottom P(B) we left it as P(B) since we already knew it. and only changed the top, which allowed us to get the probability.

OpenStudy (anonymous):

right

OpenStudy (anonymous):

You sir ( or madam) are a life saver.

OpenStudy (anonymous):

and is we put B is the probability the search is unsuccessful in region 1, how to we find \(P(B)\)? the plane had to go down somewhere in 1, 2 or 3 so we do the following \[P(B)=P(B\cap A_1)+P(B\cap A_2)+P(B\cap A_3)\] where \(A_i\) is the event the plane goes down in region \(i\) and these are computed via \[P(B\cap A_i)=P(B|A_i)P(A_i)\]

OpenStudy (anonymous):

you good for the last one?

OpenStudy (anonymous):

Don't need to do the last one. Only A and B were assigned. Prof said C was the same as B. Which makes sense. The math would look the same.

OpenStudy (anonymous):

ok sorry it took me so long i will use the excuse that the wording was confusing (which it is) but once you get the hang of bayes formula it is not so bad just keep in mind what you are really looking for

OpenStudy (anonymous):

oh yeah 3 looks identical to 2

OpenStudy (anonymous):

It's alright it took me a long time to even start asking on here.

OpenStudy (anonymous):

lol come back any time

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yw